find the frequency of the following harmonic motion model. y=0.3 cos(8t)
Answer by stanbon(75887)

find the frequency of the following harmonic motion model. y=0.3 cos(8t)
-----
Period = (2pi)/8 time/cycle
------
Frequency = 8/(2pi) cycles/time = 4/pi ~ 1.27...
--------------
Cheers,
Stan H.
-------------